OpenStudy (anonymous):

(cosx/(secx-tanx)) = 1+senx i dont know how to prove this! HELP

OpenStudy (anonymous):

I got this. Give me a sec

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\cos(x)/(\sec(x) - \tan(x))\]\[(\cos(x)/(\sec(x) - \tan(x))) * (\sec(x) + \tan(x)/(\sec(x) + \tan(x)))\]\[(\sec(x) + \tan(x)) * (\cos(x)/\sec^2(x) - \tan^2(x))\]\[(\sec(x) + \tan(x)) * \cos(x)/1\]\[\sec(x)*\cos(x) + \tan(x)*\cos(x)\]\[1 + \sin(x)\]

OpenStudy (anonymous):

That's ugly... I wish I knew better formatting tools

OpenStudy (anonymous):

cosx /(secx-tanx) = cos x / ( 1/cosx - sinx / cosx) =cosx / (1-sinx)/cosx) = cos^2x / 1-sinx = 1-sin^2x/1-sinx = (1+sinx)(1-sinx)/(1-sinx) =1+sinx

OpenStudy (anonymous):

thank you so much